answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Can you pee a marble?

Yes and no. Kidney stones feel as big as a marble. But if you swallow a marble it will come out a different way

What is pistal?

In Botany, the pistil is the female organs of a flower, comprising the stigma, style, and ovary.

The anther is the part of a stamen that contains the pollen.

What is a theoretical alternative?

A theoretical alternative refers to a concept or idea that proposes a different approach or solution to a problem, distinct from existing methods or frameworks. It often arises in academic, scientific, or philosophical discussions, where it serves as a basis for exploring new possibilities or challenging established norms. Theoretical alternatives can stimulate innovation and critical thinking, encouraging the exploration of diverse perspectives in various fields.

A distribution statement does what?

identifies who is authorized receipt of the to

How do you find the probability of drawing 5 red cards in 5 card poker hand?

A quick way is to simply count all the possible hands with red cards and compare to the total number of hands. There are 52 total cards and 26 red cards. We can use the binomial coefficient for this, otherwise known as the "choose" operator.

A = Total Hands: (52c5) = 2598960
B = Red Hands: (26c5) = 65780
C = Chance of red hand = A / B = 2.53%

Alternately you can use a permutation solution, for each card, in order, what is the chance to get a red card. There are 26 of 52 red cards on the first card, so 26/52. Then continue:

26/52 * 25/51 * 24/50 * 23/49 * 22/48 = 2.53%

If you play with the form of the equations, and understand a bit about combinations, you'll see that this second form is very closely related to the expanded version of the first form. This two methods have advantages in different situations. Here they are about the same.

What are the possible combinations of rolling two dice to get 11?

11 = 6+5 is the only solution, so there are two combinations

first dice : 6, second dice : 5

first dice : 5, second dice : 6

What is on the back side of a penny?

There is a design of the Lincoln Memorial in Washington D.C. on the back of the penny. Also are the words "United States of America" and "E Pluribus Unum" and, lastly "One Cent".

A dime and a penny are tossed and one dice is rolled what is the total number of possible outcomes?

There are 2 outcomes for the dime (H or T), 2 for the penny (H or T) and 6 for the die (1,2,3,4,5,6). In all, there are 2*2*6 = 24 outcomes. Some of them are given below in the pattern: dime, penny, die. The rest are easy to generate.

[H,H,1], [H,H,2], ... , [H,H,6],

[H,T,1], [H,T,2], ...

[T,H,1], ...

[T,T,1], ...

Why do they call tails on a coin tails?

Because there is usually the head of some emperor or king or queen on the other side. To call the opposite side "tails" just seems to be as good as calling it anything else.

How many ways can you put 5 different letters in order?

Let's start with the first.

You have 5 to choose from so that's 5.

Now for each of those 5, let's choose the second letter.

You now have 4 to choose from.

So we're up to 5 x 4.

In the third position we're down to a choice of 3, so total possibilities go up to 5x4x3.

In the fourth, we have 2 left, so the choices go up to 5x4x3x2.

After all that, there is only one left for the fifth position: just one to choose from.

So the final answer: 5x4x3x2x1.

That's called 5 factorial.

How many ways can you arrange the letters in the word THREE?

5!/2=60 [If all the letters were different, it would be 5!=120. But you need to divide by two since "e" appears twice.]

What is a zero-one law?

A zero-one law is the law that, in some mathematical circumstances, the probability of an event can only be zero or one and not any intermediate value.

Trending Questions
What is written by someone who was not present at an event called? If five dice are rolled how many possible outcomes? When did the first life forms appear on Earth? What are the odds against getting exactly two heads in three successive flips of a coin? What is the probability of rolling a sum of 7 decimal? A ball is picked at random from a box containing 4 white 3 red 11 blue and 10 green balls Find the probability of the event White or red ball is picked Give exact answer in fraction form? A bag contains 10 black marbles and 10 white marbles Five marbles are drawn and then replaced each time What is the probability of drawing five white marbles in a row? What are the most likely sums to be rolled on two regular dice? What is the anwer in getting the probability of a die less than 7? What is the probability that a five-card poker hand contains the two of diamonds and the three of spades Enter the value in decimal format and report it to four decimal places? What is the probability of having 2 red balls when chosen in random from a container has 2 red balls 4 green balls and 2 yellow balls? A 52 card deck of playing cards is shuffled and one card is dealt from the top of the deck What is the probability that it will be a black card? How do you solve Two cards are chosen at random from a deck of 52 cards without replacement What is the probability that the first card is a jack and the second card is a ten? How do you deal with cicle theorems? Which event did not happen in 1968? How did bugs rig the coin toss in bub not buddy? How do you determend probibility? What is the probability of getting at least 3 tails when tossing a coin six times? How many outcomes will there be when n coins are tossed? What is the chance you will die today?